Homework 3  EE290G Intro. to MEMS  Due 2/13/97

Reading: Kurt E. Petersen, “Silicon as a Mechanical Material”, Proc. IEEE, v. 70, no. 5, pp. 420-457, May 1982.

1. A (111) plane makes an angle of roughly 54.74 degrees with the surface of a [100] silicon wafer.

(a) Prove it.

(b) With an etch rate ratio of 400:1 between the [100] and [111] directions, the sidewalls of a pit etched in a [100] wafer will be close to the {111} planes, but not exactly parallel to them. What angle will they really make with the surface?

(c) The 400:1 ratio is possible with KOH. With EDP, a more common ratio would be 35:1 . What angle will the sidewalls make with the wafer surface in such an EDP etch?

(d) What etch rate ratio would be necessary to make 45 degree sidewalls?

2. From Petersen’s description of Tuckerman and Pease’s work on cooling, what is a lower bound on the amount of heat that could be removed from a 1 cm cube with this technique and a temperature rise of 60 degrees C? What flow rate would be required?

3. Explain how the slope of the curves in figure 52 can be used to determine the mechanical properties of the thin films. Discuss the possible sources of error?